# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v3

@@ III-B Adaptive Value Regularization on Model Data
where $T_{\widehat{\mathcal{M}}}$ represents the state transition model in the learned environment $\widehat{\mathcal{M}}$, while $T_{\mathcal{M}}$ corresponds to that in the true environment $\mathcal{M}$. Since sampling distribution for model data needs to satisfy the equation: $\sum_{s,a}\omega(s,a)=1$, the sampling distribution $\omega(s,a)$ based on finite model dataset can be denoted as:
@@
Hence, to design an accurate sampling distribution, it is crucial to compute the dynamic ratio value $T_{\widehat{\mathcal{M}}}/T_{\mathcal{M}}$ accurately. Following [28], the following equation can be derived by applying Bayes theorem:
@@
By training the discriminator networks, the gap between the learned and the actual environment can be estimated. Then, the sampling distribution of model data can be obtained. Based on this, the different sampling probabilities among model data can bring the adaptive value regularization for the model data.
@@
21:  Initialization: target network $Q_{\overline{\omega}_{1}}\leftarrow Q_{\omega_{1}}$ , $Q_{\overline{\omega}_{2}}\leftarrow Q_{\omega_{2}}$, model dataset $\mathcal{D}_{\text{model }}=\varnothing$.
32:  Use $\mathcal{D}_{\text{offline}}$ to train $N$ environment model $\{\widehat{T}_{\phi}^{i}=\mathcal{N}[\mu_{\phi}^{i}(s,a),\sum_{\phi}^{i}(s,a)]\}_{i=1}^{N}$.
@@
4:     Collect model date: rollout $(\mu(\cdot\mid s),\widehat{T}_{\phi},H)$, and add rollout date into $\mathcal{D}_{\text{model}}$ ;
5:     Update discriminator $D_{\Phi_{sas}},D_{\Phi_{sa}}$ according to Eq. (13), and calculate adaptive sampling distribution;
6:     Update policy network $\pi_{\theta}$ according to Eq. (6);
7:     Update critic network $Q_{\omega_{1}},Q_{\omega_{2}}$ according to Eq. (9);
8:     if $t$ update period = 0:
9:      Soft update periodically:
10:      $\overline{\omega}_{1,2}\leftarrow\tau\omega_{1,2}+(1-\tau)\overline{\omega}_{1,2}$;
11:     end if
12:  end for
Algorithm 1 Mildly Conservative Model-Based Offline Reinforcement Learning
III-C Algorithm Details of DOMAIN
DOMAIN employs Eq. (9) to perform mildly conservative policy evaluation, Eq. (6) for policy improvement, and Eq. (13) to train the discriminator. The term $g(s,a)$ represents the expectation of the dynamic ratio. This expectation is estimated by sampling $m$ states from a Gaussian distribution $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)=\mathcal{N}[\mu_{\phi}(s,a),\sum_{\phi}(s,a)]$ obtained from the trained model. The weight $\lambda$ of the regularization term and the rolling horizon $H$ exert a substantial influence on the agent training. The detailed analysis can be found in section V-D.
IV Theoretical Analysis of DOMAIN
In this section, we demonstrate that the adaptive sampling distribution of model data is equivalent to the adaptive adjustment of penalties for rewards. Furthermore, we prove that the learned value function in DOMAIN serves as a lower bound for the true value function in OOD region, effectively suppressing the value function in highly uncertain region. By employing a lower bound optimization strategy, DOMAIN guarantees safety policy improvement.
IV-A Adaptive Adjustment of Penalties for Rewards
@@
The detailed proof can be found in Appendix A. Theorem 1 shows that the operator $\widehat{\mathcal{T}}$ is a $\gamma$-contraction mapping in the $\mathcal{L}_{\infty}$ norm. According to the fixed point theorem, any initial $Q$ function can converge to a unique fixed point $Q^{*}$ by repeatedly applying the Bellman operator $\widehat{\mathcal{T}}^{\pi}$.
By setting the derivative of the Eq. (9) concerning $Q^{k}$ to zero during the $k$-th iteration, the following equation can be obtained:
@@
We denote the second term in Eq. (14) as $\eta(s,a)$, which serves as the adjustment of penalties for rewards. When the model data $(s,a)$ exhibits high errors, the value of the designed sampling distribution $\omega(s,a)$ increases. This indicates that the data point $(s,a)$ either resides in the OOD region or a region with a relatively low density of offline dataset, implying that $d(s,a)$ approaches zero. Consequently, $\omega(s,a)>d(s,a)$, resulting in $\eta(s,a)>0$ that serves as a penalty term. Moreover, the magnitude of penalization in the Q value increases with larger errors in the model data. In contrast, when the model data error is small, that is the data is situated in a region proximal to true environment, it is possible for $\omega(s,a)$ to be smaller than $d(s,a)$, leading to $\eta(s,a)<0$ which acts as a reward term. This encourages exploration in accurate model data region during training process.
The adaptive adjustment of reward is reflected as follows: During the value iteration process, the adjustment term for reward, $\eta(s,a)$, adapts the rewards based on the magnitude of $\omega(s,a)$ (assuming the offline data set is fixed and thus $d(s,a)$ can be considered fixed as well). This adaptive adjustment allows for penalizing model data with high errors and reduces the conservatism of the algorithm.
IV-B DOMAIN Optimizes a Low Bound Value
Considering the influence of sampling errors and model errors, we present the conditions for the existence of a lower bound on the value function in the DOMAIN, and demonstrate that DOMAIN does not underestimate the value function for all states. Additionally, we discuss a sufficient condition under which the lower bound of DOMAIN is tighter than that of COMBO, indicating that DOMAIN exhibits weaker conservatism on offline dataset.
@@ IV-B DOMAIN Optimizes a Low Bound Value
The detailed proof can be found in Appendix B. From the theorem, it can be concluded that for a sufficiently large $\lambda$ and when the aforementioned conditions are met, $V^{\pi}(s)\geq\widehat{V}^{\pi}(s)$ holds. When there is a sufficient amount of data, $\delta_{l}$ primarily stems from model errors, $C_{r,T,\delta}R_{\max}/((1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\})$ approaching zero. Therefore, when the model error is small and the model data ratio $f$ is appropriately set, a sufficiently small $\lambda$ can ensure $V^{\pi}(s)\geq\widehat{V}^{\pi}(s)$.
Furthermore, DOMAIN does not underestimate the value function for all states $s$. It can be inferred that the value function is underestimated only when the condition in Eq. (15) is satisfied. However, there is a possibility of overestimating the data with high probabilities in the distribution $d(s,a)$ of the offline data set. Since the data points $(s,a)$ with a higher frequency of occurrence in region $\mathcal{D}_{\text{offline }}$ are considered as in-distribution data, the learned environment model $\widehat{T}_{\phi}$ exhibits smaller prediction errors for such data, resulting in smaller values of $\omega(s,a)$. As a result, it becomes challenging to satisfy the condition $\sum_{a}\omega(s,a)>\xi(s)\sum_{a}d(s,a)$, leading to overestimation. This aligns with the adaptive adjustment of penalties for reward.
@@
Let $\kappa_{1}$ and $\kappa_{2}$ denote the average value functions of DOMAIN and COMBO methods respectively. Specifically, $\kappa_{1}=\mathbb{E}_{s\sim d(s),a\sim\pi(a\mid s)}[\widehat{Q}_{1}^{\pi}]$, $\kappa_{2}=\mathbb{E}_{s\sim d(s),a\sim\pi(a\mid s)}[\widehat{Q}_{2}^{\pi}]$. It holds that $\kappa_{1}>\kappa_{2}$ if the following condition is satisfied:
@@
where $\widehat{Q}_{1}^{\pi}$ and $\widehat{Q}_{2}^{\pi}$ are the $Q$-value function of DOMAIN and COMBO in learned policy $\pi$ respectively. $d(s)$ represents the state distribution of any data set, and $d_{\widehat{\mathcal{M}}}^{\pi}(s)$ represents the distribution of states under the learned model. The detailed proof of this theorem can be found in Appendix C.
When the data set corresponds to an offline dataset, the data points $(s,a)$ with higher densities in $d(s,a)$ exhibit smaller model prediction errors, implying $w(s,a)$ approaches zero. The generated model data is more likely to be concentrated in-distribution region, indicating $d_{\widehat{\mathcal{M}}}^{\pi}(s)$ is large. Hence, Eq. (16) is readily satisfied, indicating that the conservatism of DOMAIN is weaker than that of COMBO.
IV-C DOMAIN Guarantees the Safety of Policy Improvement
The goal of DOMAIN is to learn an optimal policy $\pi^{*}(a\mid s)$ in the actual MDP $\mathcal{M}$, maximizing the cumulative return $J(\mathcal{M},\pi):=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}[r(s,a)]/1-\gamma$. Building upon previous works [15, 21, 32, 33], we prove that the DOMAIN provides a $\zeta$-safe policy improvement for the behavioral policy $\pi_{b}(a\mid s)$.
@@ IV-C DOMAIN Guarantees the Safety of Policy Improvement
Figure 2: The description of Mujoco tasks, including Halfcheetah, Hopper and Walker2d tasks.
The proof of this theorem can be found in Appendix D. Here, $\mu_{2}$ represents the sampling error term, which becomes negligible when the data volume is sufficiently large. $\mu_{3}$ denotes the model error term, with a larger value indicating a larger discrepancy between the trained environment model $\widehat{T}_{\phi}$ and the true model $T$. The term $\varpi(\pi,f)=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}_{f}(s,a)}^{\pi}}[\eta(s,a)]$ measures the expected rewards and penalties under a specific policy and model data ratio, capturing the difference in the expected penalty between different policies. Therefore, $\mu_{1}$ represents the discrepancy in the expected penalties under different policies.
@@
V Experiments
This section aims to evaluate the performance of DOMAIN on standard benchmark, and gives the details of experimental setup, results and parameter analysis.
TABLE I: Base hyperparameter configuration of DOMAIN.
@@ V Experiments
V-A The Implementation Details of Experiments
Experimental Datasets. We employ Gym-Mujoco benchmark to test the performance of DOMAIN. The Gym-Mujoco tasks, including Halfcheetah-v2, Hopper-v2, and Walker2d-v2, are commonly studied in OpenAI Gym [35], simulated using the MuJoCo physics engine. Fig. 2 depicts the above tasks. The goal is to control the robot’s various joints to achieve faster and more stable locomotion. [36].
For three Mujoco tasks, three types of recorded datasets are considered: Medium, Medium-replay, and Medium-expert. The “Medium” dataset is generated by collecting 1 million samples from a policy that undergoes early stopping during online training using SAC [27]. The “Medium-replay” dataset records all samples during the training process, reaching a medium level. The “Medium-expert” dataset is a combination of expert and sub-optimal data [36].
Practical algorithm implementation details. Fig. 4 gives the network composition of the algorithm, including environment, discriminator, critic and policy network. Through the discriminator network, the model data sampling probability is calculated for the critic network update; the critic network guides the policy network update; the actions generated on the strategy network generate more model data through the environment model. Table I provides the hyper-parameters for the network model. The detailed parameter settings for each network update are given below:
@@ V-A The Implementation Details of Experiments
1) Model Training: The environment model is approximated by a probabilistic network that outputs a Gaussian distribution to obtain the next state and reward. We set $N$ as 7 and select the best 5 models. Each model consists of a 4-layer feedforward network with $200$ hidden units. The model employs maximum likelihood estimation with a learning rate of $41e-$ 4 and Adam optimizer. The $256$ data samples are randomly drawn from $\mathcal{D}_{\text{offline }}$ to train model.
@@
Figure 5: Corresponding learning curves for Table II. Each figure shows the training curve for a specific task under different datasets, where the shaded area represents the standard variance of the return.
@@
Parameter Tuning. The quality of model data is influenced by the environment model $\widehat{T}_{\phi}$, the rollout policy $\mu(\cdot\mid s)$, and the rollout horizon $H$. The weight coefficient $\lambda$ of the regularization term significantly affects the stability of agent training. The rollout policy is the current learned policy. Table III gives the results of Gym-Mujoco tasks for different parameters. The detailed parameter tuning results in different tasks are as follows.
1) Rollout horizon $H$ : The rollout horizon $H$ is chosen from the set $\{1,5\}$. We ues $H=1$ for Walker2d-v2 tasks, and $H=5$ for Hopper-v2 and Halfcheetah-v2.
@@
TABLE II: Results for D4RL tasks during the last 5 iterations of training averaged over 5 seeds. $\pm$ captures the standard deviation. Bold indicates the highest score. Highlighted numbers indicate results within 2% of the most performance algorithm.
@@
TABLE III: The results of D4RL task for different parameters. Bolded and highlighted numbers indicate the best performance
@@
V-B Experimental Results
Results on environment model prediction. This section first analyzes and visualizes the accuracy of trained dynamic model prediction before answering the above questions. Fig. 3 gives the prediction results of cumulative error in rollout lengths 1,3,5,7 for different tasks and datasets. In this figure, the predicted errors of four distinct datasets for tasks halfcheetah, hopper and walker2d are arranged from left to right. The prediction error of the trained environmental model exhibits an increasing trend with the increment of steps across all datasets. Since the model exhibits the largest prediction error in three walker2d tasks, the rollout length is set to 1 in the walker2d task to ensure the stability of the agent.
However, single-step prediction is not conducive to the expansion of the OOD data by the environment model, because most of the predicted state-action data still falls within the region. As rollout length $H$ increases, it is beneficial to explore the OOD region. Meanwhile, it is necessary to ensure the model data error is not too large. In Fig. 3, the prediction error of $H$ from 1 to 7 of the environmental model does not change too much under Halfcheetah and Hopper tasks. Within the error range, the optimal $H$ value should be determined by more experiments.
Results on the D4RL tasks. This section compares recent model-free algorithms, such as MOAC [40], IQL [37], and GORL [41], as well as model-based algorithms, including ARMOR [25], MAPLE [39], MOReL [20], COMBO [21], and RAMBO [24]. The performance of these algorithms is evaluated based on the cumulative return in the robot motion tasks within the MuJoCo environment. A higher score indicates better performance in terms of more stable and faster locomotion control. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [36]. The normalized score $\tilde{S}$ is computed by:

### v4

@@ IV-B Adaptive Value Regularization on Model Data
where $T_{\widehat{\mathcal{M}}}$ represents the state transition model in the learned environment model $\widehat{\mathcal{M}}$, while $T_{\mathcal{M}}$ corresponds to that in the true environment model $\mathcal{M}$. Since sampling distribution for model data needs to satisfy the equation: $\sum_{s,a}\omega(s,a)=1$, the sampling distribution $\omega(s,a)$ based on finite model dataset can be denoted as:
@@
Hence, to design an accurate sampling distribution, it is crucial to compute the dynamic ratio value $T_{\widehat{\mathcal{M}}}/T_{\mathcal{M}}$ precisely. Following [43], the following equation can be derived by applying Bayes’ theorem:
@@
By training the discriminator networks, the gap between the learned and actual environment can be estimated. Subsequently, the sampling distribution of model data can be obtained. Based on this, the varying sampling probabilities among the model data can introduce adaptive value regularization for the model data.
IV-C Algorithm Details of DOMAIN
DOMAIN employs Eq. (9) to perform mildly conservative policy evaluation, Eq. (6) for policy improvement, and Eq. (13) to train the discriminator. The term $g(s,a)$ represents the expectation of the dynamic ratio, which is estimated by sampling $m$ states from a Gaussian distribution $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)=\mathcal{N}[\mu_{\phi}(s,a),\sum_{\phi}(s,a)]$ obtained from the trained model. The weight $\lambda$ of the regularization term and the rolling horizon $H$ have a substantial influence on the agent training. The detailed analysis can be found in Section VI-E.
@@ IV-C Algorithm Details of DOMAIN
21:  Initialization: target network $Q_{\overline{\omega}_{1,2}}\leftarrow Q_{\omega_{1,2}}$, and model dataset $\mathcal{D}_{\text{model }}=\varnothing$.
32:  Use $\mathcal{D}_{\text{offline}}$ to train $N$ environment models $\{\widehat{T}_{\phi}^{i}\}_{i=1}^{N}$.
@@
4:     Collect model date: rollout $(\mu(\cdot\mid s),\widehat{T}_{\phi},H)$, and add rollout date (model data) into $\mathcal{D}_{\text{model}}$ ;
5:     Update discriminator $D_{\Phi_{sas}},D_{\Phi_{sa}}$ by Eq. (13), and calculate adaptive sampling distribution $\omega$;
6:     Update policy $\pi_{\theta}$ and critic $Q_{\omega_{1,2}}$ through Eq. (6) and Eq. (9) (See Table II for related parameters);
7:     if $t$ update period = 0:
8:      Soft update: $\overline{\omega}_{1,2}\leftarrow\tau\omega_{1,2}+(1-\tau)\overline{\omega}_{1,2}$;
9:     end if
10:  end for
Algorithm 1 DOMAIN: Mildly Conservative Model-Based Offline Reinforcement Learning
V Theoretical Analysis of DOMAIN
In this section, we demonstrate that the adaptive sampling distribution of model data is equivalent to the adaptive adjustment of penalties for rewards. Furthermore, we prove that the learned value function in DOMAIN serves as the lower bound for the true value in the OOD region, effectively suppressing the value under the highly uncertain region. By employing a lower-bound optimization policy, DOMAIN has a safety policy improvement guarantee.
V-A Adaptive Adjustment of Penalties for Rewards
@@
The detailed proof can be found in Appendix A (Please see Supplementary Material). Theorem 1 shows that the operator $\widehat{\mathcal{T}}$ is a $\gamma$-contraction mapping in the $\mathcal{L}_{\infty}$ norm. According to the fixed point theorem, any initial $Q$ function can converge to a unique fixed point $Q^{*}$ by repeatedly applying the Bellman operator $\widehat{\mathcal{T}}^{\pi}$.
Then, by setting the derivative of the Eq. (9) concerning $Q^{k}$ to zero during the $k$-th iteration, the following equation can be obtained:
@@
We denote the second term in Eq. (14) as $\eta(s,a)$, which serves as the adjustment of penalties for rewards. When the model data $(s,a)$ exhibits significant errors, the value of the designed sampling distribution $\omega(s,a)$ increases. This suggests that the data point $(s,a)$ either lies in the OOD region or a region with a relatively low density within the offline dataset, implying that $d(s,a)$ approaches zero. Consequently, $\omega(s,a)>d(s,a)$, resulting in $\eta(s,a)>0$ that serves as a penalty term. Moreover, the magnitude of penalization in the Q value increases with larger errors in the model data. In contrast, when the model data error is small, that is the data is situated in a region proximal to true environment, it is possible for $\omega(s,a)$ to be smaller than $d(s,a)$, leading to $\eta(s,a)<0$ which acts as a reward term. This encourages exploration in the accurate model data region during the training process.
V-B DOMAIN Optimizes a Low Bound Value
Considering the influence of sampling and model errors, we present the conditions for the existence of a lower bound on the value function in DOMAIN and demonstrate that DOMAIN does not underestimate the value function for any state. Additionally, we discuss a sufficient condition under which the lower bound of DOMAIN is tighter than that of COMBO, indicating that DOMAIN exhibits weaker conservatism.
@@ V-B DOMAIN Optimizes a Low Bound Value
The detailed proof is provided in Appendix B. Based on the theorem, for a sufficiently large $\lambda$ and when the aforementioned conditions are met, $V^{\pi}(s)\geq\widehat{V}^{\pi}(s)$ holds. when a sufficient amount of data is available, $\delta_{l}$ primarily is primarily attributed to model errors, $C_{r,T,\delta}R_{\max}/((1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\})$ approaches zero. Therefore, when the model error is small and the model data ratio $f$ is appropriately chosen, a sufficiently small $\lambda$ can ensure $V^{\pi}(s)\geq\widehat{V}^{\pi}(s)$.
Furthermore, DOMAIN does not underestimate the value function for all states $s$. The value function is underestimated only when the condition in Eq. (15) is satisfied. However, there is a possibility of overestimating the data with high probabilities in the distribution $d(s,a)$ of the offline data set. Since the data points $(s,a)$ with a higher frequency of occurrence in region $\mathcal{D}_{\text{offline }}$ are considered as in-distribution data, the learned environment model $\widehat{T}_{\phi}$ exhibits smaller prediction errors for such data, resulting in smaller values of $\omega(s,a)$. As a result, it becomes challenging to satisfy the condition $\sum_{a}\omega(s,a)>\xi(s)\sum_{a}d(s,a)$, leading to overestimation. This aligns with the adaptive adjustment of penalties for reward.
@@
Let $\kappa_{1}$ and $\kappa_{2}$ denote the average value functions of the DOMAIN and COMBO methods, respectively. Specifically, $\kappa_{1}=\mathbb{E}_{s\sim d(s),a\sim\pi(a\mid s)}[\widehat{Q}_{1}^{\pi}]$, $\kappa_{2}=\mathbb{E}_{s\sim d(s),a\sim\pi(a\mid s)}[\widehat{Q}_{2}^{\pi}]$. It holds that $\kappa_{1}>\kappa_{2}$ if the following condition is satisfied:
@@
where $\widehat{Q}_{1}^{\pi}$ and $\widehat{Q}_{2}^{\pi}$ are the $Q$-value function of DOMAIN and COMBO in learned policy $\pi$ respectively. $d(s)$ represents the state distribution of any data set, and $d_{\widehat{\mathcal{M}}}^{\pi}(s)$ represents the distribution of states under the learned model.
The detailed proof of this theorem can be found in Appendix C. When the dataset is the offline dataset, the data points $(s,a)$ with higher densities in $d(s,a)$ exhibit smaller model prediction errors, which implies $w(s,a)$ approaches zero. The generated model data is more likely to be concentrated in-distribution region, indicating $d_{\widehat{\mathcal{M}}}^{\pi}(s)$ is large. Hence, Eq. (16) is readily satisfied, indicating that the conservatism of DOMAIN is weaker than that of COMBO.
V-C DOMAIN Guarantees the Safety of Policy Improvement
The offline dataset is the only dataset available for policy learning of DOMAIN, and the policy in the offline dataset can be described as behavior policy $\pi_{b}(a\mid s)$. The goal of DOMAIN is to learn an optimal policy $\pi^{*}(a\mid s)$ in the actual MDP $\mathcal{M}$ through maximizing the cumulative return $J(\mathcal{M},\pi):=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}[r(s,a)]/1-\gamma$, and the performance gap between agent under optimal policy and behavioral policy should be bounded. Building upon previous works [22, 28, 47, 48], we prove that the DOMAIN provides a $\zeta$-safe policy improvement for the behavioral policy.
@@ V-C DOMAIN Guarantees the Safety of Policy Improvement
Figure 2: The description of Mujoco tasks, including Halfcheetah, Hopper, and Walker2d tasks.
The proof of this theorem can be found in Appendix D. Here, $\mu_{2}$ represents the sampling error term, which becomes negligible when the data volume is sufficiently large. The term $\mu_{3}$ denotes the model error term, where a larger value indicates a greater discrepancy between the trained environment model $\widehat{T}_{\phi}$ and the true model $T$. Additionally, $\varpi(\pi,f)=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}_{f}(s,a)}^{\pi}}[\eta(s,a)]$ quantifies the expected rewards and penalties under a specific policy and model data ratio, capturing the difference in the expected penalty between different policies. Therefore, $\mu_{1}$ represents the discrepancy in the expected penalties under different policies.
@@
VI Experiments
This section aims to evaluate the performance of DOMAIN on standard benchmarks, and gives the details of experimental setup, results on D4RL tasks, computation complexity, robust performance, and key parameters analysis.
TABLE II: Base hyperparameter configuration of DOMAIN.
@@ VI Experiments
VI-A The Implementation Details of Experiments
Experimental Datasets. We employ Gym-Mujoco tasks to test the performance of DOMAIN. The Gym-Mujoco tasks, such as Halfcheetah-v2, Hopper-v2, and Walker2d-v2, are included in the D4RL benchmark [50], which are widely used in offline RL. Fig. 2 depicts the above tasks. The goal is to control the robot’s various joints to achieve faster and more stable locomotion. [51].
In our paper, we select $9$ datasets covering $3$ tasks. For three Gym-Mujoco tasks, three types of recorded datasets are considered: Medium, Medium-replay, and Medium-expert. The “Medium” dataset is generated by collecting 1 million samples from a policy that undergoes early stopping during online training using SAC [42]. The “Medium-replay” dataset records all samples during the training process, reaching a medium level. The “Medium-expert” dataset is a combination of expert and sub-optimal data with a total 2 million samples [51].
Practical algorithm implementation details. Fig. 4 gives the network composition of the algorithm, including environment, discriminator, critic, and policy network. Through the discriminator network, the model data sampling probability is calculated for the critic network update; the critic network guides the policy network update; the actions generated on the strategy network generate more model data through the environment model; the initial state $s$ is randomly sampled from the offline dataset and used to the generation of model data. Table II provides the hyperparameters for the network model. The detailed parameter settings are given below:
@@ VI-A The Implementation Details of Experiments
1) Model Training: The environment model is approximated by a probabilistic network that outputs a Gaussian distribution to obtain the next state and reward. We set $N$ as 7 and select the best 5 models. Each model consists of a 4-layer feedforward network with $200$ hidden units. The model employs maximum likelihood estimation with a learning rate of $41e-$ 4 and Adam optimizer. The $256$ data samples are randomly drawn from $\mathcal{D}_{\text{offline }}$ to train the model.
@@
Figure 5: Corresponding learning curves for Table III. Each figure shows the training curve for a specific task under different datasets, where the shaded area represents the standard variance of the return.
@@
Parameter Tuning. The quality of model data is influenced by the environment model $\widehat{T}_{\phi}$, the rollout policy $\mu(\cdot\mid s)$, and the rollout horizon $H$. The weight coefficient $\lambda$ of the regularization term significantly affects the stability of agent training. The rollout policy is the current learned policy. The parameters $H$ and $\lambda$ are two critical hyperparameters that require meticulous tuning to ensure optimal performance. The selection of the parameter $H$ is closely tied to the accuracy of the training environment model, and the regularization approach employed by the algorithm to handle erroneous model data. The choice of the parameter $\lambda$ is determined by the inherent distributional characteristics of the dataset. The detailed parameter analysis can be found in section VI-E.
Moreover, note that RL is characterized by high training costs and instability, with the ideal approach being the identification of a set of hyperparameters that generalize well across various tasks, rather than requiring frequent tuning for each task. In practical applications, the computational cost of exhaustive parameter search is prohibitively high. This study is based on the COMBO framework, where the selection of parameters has two options. To maintain experimental rigor and comparative fairness, the set of parameter candidates is the same as COMBO. Table IV gives the results of Gym-Mujoco tasks under different parameters. Based on this table, parameters $H$ and $\lambda$ for different tasks are set as:
1) Rollout horizon $H$ : The rollout horizon $H$ is chosen from the set $\{1,5\}$. We uses $H=1$ for Walker2d-v2 tasks, and $H=5$ for Hopper-v2 and Halfcheetah-v2.
@@
TABLE III: Results for D4RL tasks during the last 5 iterations of training averaged over 5 seeds. $\pm$ captures the standard deviation. Bold indicates the highest score. Highlighted numbers indicate results within 2% of the most performing algorithm.
@@
TABLE IV: The results of D4RL task for different parameters. Bolded and highlighted numbers indicate the best performance.
@@
VI-B Experimental Results
Results on environment model prediction. This section first analyzes and visualizes the accuracy of the trained environment model prediction before answering the above questions. Fig. 3 gives the prediction results of cumulative error in rollout lengths 1,3,5, and 7 for different tasks and datasets, respectively. In this figure, the predicted errors of four distinct datasets for tasks Halfcheetah, Hoppe, and Walker2d are arranged from left to right. The prediction error of the trained environmental model exhibits an increasing trend with the increment of steps across all datasets. Since the model exhibits the largest prediction error of the three Walker2d tasks, the rollout length is set to 1 in the Walker2d task to ensure the agent’s stability.
However, single-step prediction is not conducive to the expansion of the OOD data by the environment model, because most of the predicted state-action data still falls within the region. As the rollout length $H$ increases, it is beneficial to explore the OOD region. Meanwhile, it is necessary to ensure the model data error is not too large. In Fig. 3, the prediction error of $H$ from 1 to 7 of the environmental model does not change too much under Halfcheetah and Hopper tasks. Within the error range, the optimal $H$ value should be determined by more experiments.
Results on the D4RL tasks. This section compares recent model-free algorithms, such as MOAC [29], IQL [52], and GORL [30], as well as model-based algorithms, including ARMOR [37], MAPLE [31], MOReL [27], COMBO [28], and RAMBO [36]. The results of the above methods are all based on the D4RL benchmark. The performance of these algorithms is evaluated based on the cumulative return in the robot motion tasks within the MuJoCo environment. A higher score indicates better performance in terms of more stable and faster locomotion control. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [51]. The normalized score $\tilde{S}$ is computed by: